# Two late-T dwarfs at kiloparsec distances revealed by JWST UNCOVER survey

### v1

@@ Abstract
The launch of the James Webb Space Telescope (JWST) has ushered in a new era of BD research. Its Near Infrared Camera (NIRCam; Rieke et al., 2023) provides photometric data from 0.6 to 5.0 m with unprecedented depth, reaching AB magnitudes of 30. This capability enables the detection of distant BDs beyond the Galactic plane. Several studies have predicted the number of BDs observable in JWST deep surveys (Ryan and Reid, 2016; Aganze et al., 2022; Casey et al., 2023), and multiple BD candidates have been identified in JWST deep multi-band imaging surveys (Glazebrook et al., 2023; Nonino et al., 2023; Wang et al., 2023b; Hainline et al., 2024b; Holwerda et al., 2024; Chen et al., 2025; Hainline et al., 2025). Using its Near Infrared Spectrograph (NIRSpec; Jakobsen et al., 2022), several of these BD candidates in JWST deep fields have been spectroscopically confirmed (Langeroodi and Hjorth, 2023; Burgasser et al., 2024; Hainline et al., 2024a, 2025; Tu et al., 2025).
We aimed to identify new BD candidates through spectral energy distribution (SED) model fitting using data from the JWST Cycle 1 treasury program Ultradeep NIRSpec and NIRCam ObserVations before the Epoch of Reionization (UNCOVER; Bezanson et al., 2024) and the Cycle 2 MegaScience Survey (Suess et al., 2024). In this paper, we present two robust late T dwarf candidates and one possible L subdwarfs identified in the Abell 2744 field. Section 2 describes our selection method for BD candidates. Section 3 presents the properties of the new BD candidates. In Section 4, we discuss the Galactic distribution of UNCOVER BDs. Finally, we summarize our findings in Section 5.
@@ 2.2 Initial selection by colours
Both models and observed spectra (Marley et al., 2021; Beiler et al., 2024) demonstrate that T dwarfs exhibit suppressed flux in the F277W band due to water and methane absorption features, as well as collision-induced absorption from H2–H2 (CIA H2; e.g., Borysow et al., 1997). Consequently, T dwarfs display blue F115WF277W and F150WF277W colours, alongside red F277WF444W colours, as reflected in their spectral energy distributions (e.g., Wang et al., 2023b; Hainline et al., 2024b). Figure 1 presents UNCOVER sources in F115WF277W versus F277WF444W and F150WF277W versus F277WF444W colour–colour spaces, together with theoretical colours of ultracool dwarfs from the Sonora Bobcat models (Marley et al., 2021).
@@ 2.3 SED fitting to stellar atmospheric models
To determine the atmospheric properties of the 768 colour-selected BD candidates, we constructed their SEDs using photometric data from UNCOVER and fitted their observed SEDs to various atmospheric models using the Virtual Observatory SED Analyzer (VOSA; Bayo et al., 2008; Rodrigo et al., 2024). Our SED fitting process employed different sets of atmospheric models available through VOSA: BT-Settl (Allard et al., 2012; Asplund et al., 2009; Caffau et al., 2011), ATMO (Phillips et al., 2020), Morley et al. (2012, 2014), and Saumon et al. (2012). BT-Settl models provide three parameters: , surface gravity (), and metallicity [M/H], while the other models are limited to solar metallicity. We selected only sources with more than three photometric data points and signal-to-noise ratio (SNR) 3 for fitting. The final result was determined by selecting the model that yielded the lowest value.
@@
- Notes.
@@
Notes.
aBased on correlations between spectral types and F444W-band absolute magnitudes (Burgasser et al., 2024). bGalactic height. cClassified by NIRSpec spectra (Burgasser et al., 2024). dDerived by F277W F444W versus spectral type correlation in Fig. 6-b. eDerived by spectral type – correlation for L subdwarfs (Zhang et al., 2018).
@@ 3.2 A possible L subdwarf
UNCOVER-DR3-59579 shows a significantly better fit to stellar models than to galaxy templates (Fig. 4, bottom-right panel). Its SED displays a V-shaped profile similar to that of UNCOVER-BD-1, with a flux minimum at F277W. Although the absence of F410M photometry prevents full characterisation of the 4 m peak, the object exhibits clear detections across most NIRCam bands, supporting its classification as a BD candidate. Note that the F090M band flux is not included in the SED fitting because it appears undetected in the F090W band image.
UNCOVER-DR3-59579 shares notable similarities with UNCOVER-BD-1, the T1 subdwarf with = 1300 K and [M/H] = beyond the thick disk (Burgasser et al., 2024). The best-fitting solar-metallicity ATMO model yields = 1600 K for UNCOVER-DR3-59579, corresponding to a spectral type of L5.1 for field L dwarfs. However, its Galactic height ( pc; Section 4.1) beyond the thick disc and spectral similarities to UNCOVER-BD-1 suggest it may possess subsolar metallicity. Consequently, its true could be cooler than indicated by solar-metallicity models, potentially placing it in the late-L or early-T subdwarf regime.
@@ 3.3 SED fitting to low metallicity models
The best-fitting LOWZ model ( = 600 K, log = 5.25, [M/H] = 0.0) suggests solar metallicity for UNCOVER-BD-4, which is slightly warmer than UNCOVER-BD-3. The fitting result for UNCOVER-BD-4 should be more reliable than that for UNCOVER-BD-3 owing to its brighter magnitudes and the additional F090W band measurement. The derived of 600 K corresponds to a spectral type of T8.7 according to the spectral type– relation (Kirkpatrick et al., 2021). This is slightly lower than that indicated by the F277WF444W versus correlation (Fig. 6a). The same discrepancy is observed for UNCOVER-BD-3. The F277WF444W versus spectral type correlation in Fig. 6b suggests that UNCOVER-BD-4 has a spectral type of T8, as it exhibits a slightly bluer F277WF444W colour than the T8–T9 type UNCOVER-BD-3. Since F277WF444W colour is a more direct observable than , we adopt this spectral type estimate.
UNCOVER-BD-5 also benefits from full F090W to F480M band coverage. The best-fitting LOWZ model ( = 750 K, log = 5.25, [M/H] = ) indicates subsolar metallicity for UNCOVER-BD-5. This metallicity is consistent with thick disc membership, which is further supported by its Galactic height ( pc; see Section 4.1). A of 750 K corresponds to a spectral type of T7.5 for field BDs. The F277WF444W versus spectral type correlation in Fig. 6b suggests a T7 classification for UNCOVER-BD-5. Considering the subsolar metallicity inferred from its best-fitting model, we classify UNCOVER-BD-5 as a T7 subdwarf.
@@ 4.1 Photometric distance
Figure 6c demonstrates that the F277WF444W colour is well correlated with the F444W-band absolute magnitude (). We therefore used the F277WF444W colours of these UNCOVER BDs to derive their values, and subsequently estimated photometric distances by combining their observed F444W-band magnitudes with the corresponding to their F277WF444W colours. These photometric distances are indicated in Fig. 6c and listed in Table 3.
@@ 4.2 Survey area and depth
To estimate the surface and space density of BDs in UNCOVER, we computed the footprint of the common area covered by the F277W, F444W, and F115W/F150W band mosaic images in DR3 of the UNCOVER and MegaScience surveys. Coverage in these bands is necessary for our initial colour selection (Fig. 1). We quantified the footprint of the DR3 mosaic images via FITS header metadata and pixel data properties (Bezanson et al., 2024; Suess et al., 2024). The per-pixel solid angular area was calculated from the determinant of the WCS CD matrix elements in the FITS primary header, which describe the linear transformation from pixel coordinates to celestial coordinates. As the UNCOVER observations are arranged in a mosaic containing some unobserved/blank pixels, we masked the imaging array to retain only valid observed pixels. We multiplied the number of valid pixels by the single-pixel angular area, obtaining a total effective sky coverage of 53.4 arcmin2.
Reliable identification of photometric candidates of T dwarfs in UNCOVER requires robust detections across multiple NIRCam bands, particularly in F277W where T dwarfs are faintest. Consequently, we define the survey depth using the F277W magnitude of the faintest object in our sample (UNCOVER-BD-3 with F277W = 29.24 AB mag) rather than the nominal UNCOVER limiting magnitudes. Fig. 6d shows the F277WF444W versus correlation, with UNCOVER BDs overplotted using F277W-band magnitudes adjusted for their respective distances.
@@ 4.3 Surface and space number densities
Table 4 compares the space density of T5-T9 dwarfs within the UNCOVER detection cone – including UNCOVER-BD-2 and UNCOVER-BD-3 from Burgasser et al. (2024) – with the local population within 20 pc (Kirkpatrick et al., 2021), as a function of and spectral type. The number density of T5–T7 dwarfs in the UNCOVER cone extending to 2.32.0 kpc is substantially lower than in the 20 pc sample, with only UNCOVER-BD-2 representing this spectral range. Similarly, the T7 dwarf density out to 1.5 kpc falls significantly below solar neighbourhood values, represented solely by UNCOVER-BD-5. In contrast, the T8T8.9 dwarf density appears comparable to local values. This suggests that the number density of T8-T8.9 dwarfs remains relatively constant throughout the thick disk. This trend may partially reflect the extended cooling timescales of thick disc BDs, most of which have evolved to late-T spectral types (e.g., fig. 5 in Zhang et al., 2019).
@@ 5 Summary and conclusions
In the UNCOVER field, we determine a T dwarf surface density of 0.094 arcmin-2 to a limiting depth of F115W = 28.03 mag, lower than previous predictions but consistent at the order-of-magnitude level. We also calculated BD space number densities within the UNCOVER detection cone. The number density of T5–T6.9 dwarfs extending to 2.21.6 kpc falls substantially below solar neighbourhood values. The T7 dwarf density out to 1.1 kpc, while slightly higher, remains significantly lower than local measurement. Conversely, the number density of T8–T8.9 dwarfs within the thick disk appears comparable to solar neighbourhood values.
This study demonstrates the capability of JWST deep surveys to identify distant, cool substellar objects. Spectroscopic confirmation and characterisation of these candidates will refine distance constraints and provide crucial insights into the Galactic distribution and properties of BDs, advancing our understanding of substellar formation and evolutionary processes.
@@ Acknowledgements
This work is based in part on observations made with the NASA/ESA/CSA James Webb Space Telescope (JWST). Some of the data presented in this paper were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute. These observations are associated with JWST programs JWST-GO-2561, JWST-GO-2883, JWST-GO-3516, JWST-GO-3538, JWST-GO-4111, JWST-ERS-1324, and JWST-DD-2767. This publication makes use of VOSA, developed under the Spanish Virtual Observatory project supported from the Spanish MICINN through grant AyA2008-02156. MCGO acknowledges financial support from the Agencia Estatal de Investigación (AEI/10.13039/501100011033) of the Ministerio de Ciencia e Innovación and the ERDF “A way of making Europe” through project PID2022-137241NB-C42. The authors thank the referee for the useful and constructive comments.
@@ Appendix A High-redshift galaxies and ambiguous objects
Table 5 presents four spectroscopically confirmed high-redshift galaxies and one strong photometric candidate galaxy (UNCOVER-DR3-68341), as well as four ambiguous objects (UNCOVER-DR3-72947, UNCOVER-DR3-3847, UNCOVER-DR3-13700, and UNCOVER-DR3-15552) for which the photometry cannot reliably distinguish between high-redshift galaxies and brown dwarfs. Notably, the SEDs of the confirmed galaxies show better fits to substellar atmosphere models than to galaxy templates (Table 5), highlighting the limitations of BD identification based solely on NIRCam photometry. The NIRSpec spectra of the four confirmed galaxies are displayed in Fig. 9, while the SEDs of the other five objects are shown in Fig. 10.

### v2

@@ Abstract
The launch of the James Webb Space Telescope (JWST) has ushered in a new era of BD research. Its Near Infrared Camera (NIRCam; Rieke et al., 2023) provides photometric data from 0.6 to 5.0 m with unprecedented depth, reaching AB magnitudes of 30. This capability enables the detection of distant BDs beyond the Galactic plane. Several studies have predicted the number of BDs observable in JWST deep surveys (Ryan and Reid, 2016; Aganze et al., 2022; Casey et al., 2023), and multiple BD candidates have been identified in JWST deep multi-band imaging surveys (Glazebrook et al., 2023; Nonino et al., 2023; Wang et al., 2023b; Hainline et al., 2024b, 2025; Holwerda et al., 2024; Chen et al., 2025). its Near Infrared Spectrograph (NIRSpec; Jakobsen et al., 2022), several of these BD candidates in JWST deep fields have been spectroscopically confirmed (Langeroodi and Hjorth, 2023; Burgasser et al., 2024; Hainline et al., 2024a; Tu et al., 2025a, b; Morrissey et al., 2025).
We aimed to identify new BD candidates through spectral energy distribution (SED) model fitting using data from the JWST Cycle 1 treasury programme Ultradeep NIRSpec and NIRCam ObserVations before the Epoch of Reionization (UNCOVER; Bezanson et al., 2024) and the Cycle 2 MegaScience Survey (Suess et al., 2024). In this paper, we present two robust late T dwarf candidates and one possible L subdwarfs identified in the Abell 2744 field. Section 2 describes our selection method for BD candidates. Section 3 presents the properties of the new BD candidates. In Section 4, we discuss the Galactic distribution of UNCOVER BDs. Finally, we summarize our findings in Section 5.
@@ 2.2 Initial selection by colours
Both models and observed spectra (Marley et al., 2021; Beiler et al., 2024) demonstrate that T dwarfs exhibit suppressed flux in the F277W band due to water and methane absorption features, as well as collision-induced absorption from H2–H2 (CIA H2; e.g., Borysow et al., 1997). Consequently, T dwarfs display blue F115WF277W and F150WF277W colours, alongside red F277WF444W colours, as reflected in their SEDs (e.g., Wang et al., 2023b; Hainline et al., 2024b). Figure 1 presents UNCOVER sources in F115WF277W versus F277WF444W and F150WF277W versus F277WF444W colour–colour spaces, together with theoretical colours of ultracool dwarfs from the Sonora Bobcat models (Marley et al., 2021).
@@ 2.3 SED fitting to stellar atmospheric models
To determine the atmospheric properties of the 768 colour-selected BD candidates, we constructed their SEDs using photometric data from UNCOVER and fitted their observed SEDs to various atmospheric models using the Virtual Observatory SED Analyzer (VOSA; Bayo et al., 2008; Rodrigo et al., 2024). Our SED fitting process employed different sets of atmospheric models available through VOSA: BT-Settl (Allard et al., 2012; Asplund et al., 2009; Caffau et al., 2011), ATMO (Phillips et al., 2020), Morley et al. (2012, 2014), and Saumon et al. (2012). BT-Settl models provide three parameters: , surface gravity (), and metallicity [M/H], while the other models are limited to solar metallicity. We selected only sources with more than three photometric data points and SNR 3 for fitting. The final result was determined by selecting the model that yielded the lowest value.
@@
bKnown BDs confirmed by NIRSpec spectroscopy (Burgasser et al., 2024).
@@
aBased on correlations between spectral types and F444W-band absolute magnitudes (Burgasser et al., 2024).
bGalactic height.
cClassified by NIRSpec spectra (Burgasser et al., 2024).
dDerived by F277W F444W versus spectral type correlation in Fig. 6(b).
eDerived by spectral type – correlation for L subdwarfs (Zhang et al., 2018).
@@ 3.2 A possible L subdwarf
UNCOVER-DR3-59579 shows a significantly better fit to stellar models than to galaxy templates (Fig. 4, bottom-right panel). Its SED displays a V-shaped profile similar to that of UNCOVER-BD-1, with a flux minimum at F277W. Although the absence of F410M photometry prevents full characterization of the 4 m peak, the object exhibits clear detections across most NIRCam bands, supporting its classification as a BD candidate. Note that the F090M band flux is not included in the SED fitting because it appears undetected in the F090W band image.
UNCOVER-DR3-59579 shares notable similarities with UNCOVER-BD-1, the T1 subdwarf with = 1300 K and [M/H] = beyond the thick disc (Burgasser et al., 2024). The best-fitting solar-metallicity ATMO model yields = 1600 K for UNCOVER-DR3-59579, corresponding to a spectral type of L5.1 for field L dwarfs. However, its Galactic height ( pc; Section 4.1) beyond the thick disc and spectral similarities to UNCOVER-BD-1 suggest it may possess subsolar metallicity. Consequently, its true could be cooler than indicated by solar-metallicity models, potentially placing it in the late-L or early-T subdwarf regime.
@@ 3.3 SED fitting to low metallicity models
The best-fitting LOWZ model ( = 600 K, log = 5.25, [M/H] = 0.0) suggests solar metallicity for UNCOVER-BD-4, which is slightly warmer than UNCOVER-BD-3. The fitting result for UNCOVER-BD-4 should be more reliable than that for UNCOVER-BD-3 owing to its brighter magnitudes and the additional F090W band measurement. The derived of 600 K corresponds to a spectral type of T8.7 according to the spectral type– relation (Kirkpatrick et al., 2021). This is slightly lower than that indicated by the F277WF444W versus correlation (Fig. 6(a)). The same discrepancy is observed for UNCOVER-BD-3. The F277WF444W versus spectral type correlation in Fig. 6(b) suggests that UNCOVER-BD-4 has a spectral type of T8, as it exhibits a slightly bluer F277WF444W colour than the T8–T9 type UNCOVER-BD-3. Since F277WF444W colour is a more direct observable than , we adopt this spectral type estimate.
UNCOVER-BD-5 also benefits from full F090W to F480M band coverage. The best-fitting LOWZ model ( = 750 K, log = 5.25, [M/H] = ) indicates subsolar metallicity for UNCOVER-BD-5. This metallicity is consistent with thick disc membership, which is further supported by its Galactic height ( pc; see Section 4.1). A of 750 K corresponds to a spectral type of T7.5 for field BDs. The F277WF444W versus spectral type correlation in Fig. 6(b) suggests a T7 classification for UNCOVER-BD-5. Considering the subsolar metallicity inferred from its best-fitting model, we classify UNCOVER-BD-5 as a T7 subdwarf.
@@ 4.1 Photometric distance
Figure 6(c) demonstrates that the F277WF444W colour is well correlated with the F444W-band absolute magnitude (). We therefore used the F277WF444W colours of these UNCOVER BDs to derive their values, and subsequently estimated photometric distances by combining their observed F444W-band magnitudes with the corresponding to their F277WF444W colours. These photometric distances are indicated in Fig. 6(c) and listed in Table 3.
@@ 4.2 Survey area and depth
To estimate the surface and space density of BDs in UNCOVER, we computed the footprint of the common area covered by the F277W, F444W, and F115W/F150W band mosaic images in DR3 of the UNCOVER and MegaScience surveys. Coverage in these bands is necessary for our initial colour selection (Fig. 1). We quantified the footprint of the DR3 mosaic images via FITS header metadata and pixel data properties (Bezanson et al., 2024; Suess et al., 2024). The per-pixel solid angular area was calculated from the determinant of the WCS CD matrix elements in the FITS primary header, which describe the linear transformation from pixel coordinates to celestial coordinates. As the UNCOVER observations are arranged in a mosaic containing some unobserved/blank pixels, we masked the imaging array to retain only valid observed pixels. We multiplied the number of valid pixels by the single-pixel angular area, obtaining a total effective sky coverage of 53.4 arcmin2 (Fig. 8).
Reliable identification of photometric candidates of T dwarfs in UNCOVER requires robust detections across multiple NIRCam bands, particularly in F277W where T dwarfs are faintest. Consequently, we define the survey depth using the F277W magnitude of the faintest object in our sample (UNCOVER-BD-3 with F277W = 29.24 AB mag) rather than the nominal UNCOVER limiting magnitudes. Fig. 6(d) shows the F277WF444W versus correlation, with UNCOVER BDs overplotted using F277W-band magnitudes adjusted for their respective distances.
@@ 4.3 Surface and space number densities
Table 4 compares the space density of T5-T9 dwarfs within the UNCOVER detection cone – including UNCOVER-BD-2 and UNCOVER-BD-3 from Burgasser et al. (2024) – with the local population within 20 pc (Kirkpatrick et al., 2021), as a function of and spectral type. The number density of T5–T7 dwarfs in the UNCOVER cone extending to 2.32.0 kpc is substantially lower than in the 20 pc sample, with only UNCOVER-BD-2 representing this spectral range. Similarly, the T7 dwarf density out to 1.5 kpc falls significantly below solar neighbourhood values, represented solely by UNCOVER-BD-5. In contrast, the T8T8.9 dwarf density appears comparable to local values. This suggests that the number density of T8-T8.9 dwarfs remains relatively constant throughout the thick disc. This trend may partially reflect the extended cooling timescales of thick disc BDs, most of which have evolved to late-T spectral types (e.g., fig. 5 in Zhang et al., 2019).
@@ 5 Summary and conclusions
In the UNCOVER field, we determine a T dwarf surface density of 0.094 arcmin-2 to a limiting depth of F115W = 28.03 mag, lower than previous predictions but consistent at the order-of-magnitude level. We also calculated BD space number densities within the UNCOVER detection cone. The number density of T5–T6.9 dwarfs extending to 2.21.6 kpc falls substantially below solar neighbourhood values. The T7 dwarf density out to 1.1 kpc, while slightly higher, remains significantly lower than local measurement. Conversely, the number density of T8–T8.9 dwarfs within the thick disc appears comparable to solar neighbourhood values.
This study demonstrates the capability of JWST deep surveys to identify distant, cool substellar objects. Spectroscopic confirmation and characterization of these candidates will refine distance constraints and provide crucial insights into the Galactic distribution and properties of BDs, advancing our understanding of substellar formation and evolutionary processes.
@@ Acknowledgements
This work is based on observations made with the NASA/ESA/CSA JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs JWST-GO-2561, JWST-GO-2883, JWST-GO-3516, JWST-GO-3538, JWST-GO-4111, JWST-ERS-1324, and JWST-DD-2767. This publication makes use of VOSA, developed under the Spanish Virtual Observatory project supported from the Spanish MICINN through grant AyA2008-02156. MCGO acknowledges financial support from the Agencia Estatal de Investigación (AEI/10.13039/501100011033) of the Ministerio de Ciencia e Innovación and the ERDF “A way of making Europe” through project PID2022-137241NB-C42. The authors thank the referee for the useful and constructive comments.
@@ Appendix A High-redshift galaxies and ambiguous objects
Table 5 presents four spectroscopically confirmed high-redshift galaxies and one strong photometric candidate galaxy (UNCOVER-DR3-68341), as well as four ambiguous objects (UNCOVER-DR3-72947, UNCOVER-DR3-3847, UNCOVER-DR3-13700, and UNCOVER-DR3-15552) for which the photometry cannot reliably distinguish between high-redshift galaxies and BDs. Notably, the SEDs of the confirmed galaxies show better fits to substellar atmosphere models than to galaxy templates (Table 5), highlighting the limitations of BD identification based solely on NIRCam photometry. The NIRSpec spectra of the four confirmed galaxies are displayed in Fig. 9, while the SEDs of the other five objects are shown in Fig. 10.